Lisa's and her dads ages combined are 51. Her dads age is 4 less than 4 times her age How old is Lisa?

Answer by macston(5194)

L=Lisa; D=dad
D=4L-4
L+D=51 Substitute for D
L+4L-4=51
5L=55
L=11 ANSWER Lisa is 11 years old.
CHECK:
L+D=51yrs
D=51yrs-L
D=51yrs-11yrs=40 yrs Dad is 40 yrs old.
40yrs=4(11yrs)-4yrs
40yrs=44yrs-4yrs
40yrs=40yrs